OpenStudy (anonymous):

What's an equation for A new plasma-screen television costs $5250. A family makes a down payment of $552 and pays off the balance in 24 equal monthly payments.

OpenStudy (anonymous):

No answer. Just equation

OpenStudy (lilg132):

C = D - C / M Cost = Deposit - Cost / Months 5250 = 5250 - 552 / 24

OpenStudy (anonymous):

Thank you. :)

OpenStudy (anonymous):

I would rather write it more clearly as: Monthly payment = (5250-552)/24

OpenStudy (lilg132):

Monthly payments = Cost - Deposit / Months

OpenStudy (anonymous):

Yes, but arranged more clearly: Monthly payments = (cost - deposit)/months